OpenStudy (anonymous):

Form the composite function (S ◦ R)(p). R=p-1200 S=0.92p and the composite function of (R ◦ S)(p) is 0.92p-1200

OpenStudy (freckles):

\[(S ^o R)(p)=S(R(p))\] first replace the inner function R(p) there with what it equals \[S(p-1200)\] now find S(p-1200)

OpenStudy (freckles):

that is wherever you see a p in S=0.92p you will write instead of p p-1200

OpenStudy (freckles):

for example S(fish+cat)=0.92(fish+cat)

OpenStudy (anonymous):

I still don't get it

OpenStudy (freckles):

\[S(p)=0.92p\]

OpenStudy (freckles):

replace the p's with (p-1200)'s

OpenStudy (freckles):

to find S(p-1200)

OpenStudy (anonymous):

1. 1.08p − 1200 2. None of these 3. 0.92p − 1200 4. 1.08p − 1104 5. p − 1104 6. p − 1200 7. 0.92p − 1104 these are the answer choices and What I think your trying to point out is not there

OpenStudy (freckles):

i see it to get their answer you have to distribute

OpenStudy (freckles):

a(b-c)=ab-ac

OpenStudy (freckles):

have you found S(p-1200) yet?

OpenStudy (anonymous):

is it 7

OpenStudy (freckles):

yes \[S(R(p))=S(p-1200)=.92(p-1200)=.92p-1104\]

OpenStudy (anonymous):

I get it Now thank you